OpenStudy (anonymous):

Which equation is in point-slope form for the given point and slope? Point (1,-7);slope-2/3 A- y-7=2/3(x-1) B- y+7=2/3(x+1) C- y+7=-2/3(x-1) D- y-7=-2/3(x+1)

iYuko (iyuko):

point-slope form is: y - y1 = m(x - x1)

iYuko (iyuko):

M is the slope

iYuko (iyuko):

So just plug them in

iYuko (iyuko):

y + 7 = -2/3 (x - 1)

iYuko (iyuko):

\[c\]

OpenStudy (anonymous):

Thank you